OpenStudy (blackstreet23):

Use implicit differentiation to find the specified derivative. y = sin x, dx/dy

OpenStudy (perl):

There are two ways to do this. First approach: You can take the implicit derivative with respect to x, then flip it to dx/dy. $$ \Large{ \frac{d}{dx} (y) = \frac{d}{dx} (\sin x ) \\ \iff \\ \frac{dy}{dx} = \cos x \\ \iff \\ \frac{dx}{dy} = \frac{1}{\cos x } \\ \therefore \\ \frac{dx}{dy} = \sec x } $$ Second approach Take the implicit derivative with respect to y $$ \Large{ \frac{d}{dy} (y) = \frac{d}{dy} (\sin x ) \\ \iff \\ 1 = \cos x \cdot \frac{dx}{dy} \\ \iff \\ \frac{dx}{dy} = \frac{1}{\cos x } \\ \therefore \\ \frac{dx}{dy} = \sec x } $$